OpenStudy (anonymous):

integral of arcsinx/squareroot(1-x^2)

OpenStudy (lone001):

\[\int\limits\limits\limits_{}^{}arcsinx*dx/\sqrt{1-x ^{2}}=\left[ dx/\sqrt{1-x ^{2}}=d(arcsinx) \right]=\int\limits\limits\limits_{}^{}arcsinxd(arcsinx)=\arcsin ^{2}x/2\]